# EliGen: Entity-Level Controlled Image Generation with Regional Attention

### v1

@@ Abstract
Recent advancements in diffusion models have significantly advanced text-to-image generation, yet global text prompts alone remain insufficient for achieving fine-grained control over individual entities within an image. To address this limitation, we present EliGen, a novel framework for Entity-Level controlled Image Generation. We introduce regional attention, a mechanism for diffusion transformers that requires no additional parameters, seamlessly integrating entity prompts and arbitrary-shaped spatial masks. By contributing a high-quality dataset with fine-grained spatial and semantic entity-level annotations, we train EliGen to achieve robust and accurate entity-level manipulation, surpassing existing methods in both positional control precision and image quality. Additionally, we propose an inpainting fusion pipeline, extending EliGen’s capabilities to multi-entity image inpainting tasks. We further demonstrate EliGen’s flexibility by integrating it with community models such as IP-Adapter and MLLM, unlocking new creative possibilities. The source code, model, and dataset are published at the Project Page.
Figure 1: EliGen enables spatial and semantic control of each entity. (a) By incorporating local prompts and masks for each entity, it generates images with specific layouts and details. Unlike previous models restricted to rectangular controls, EliGen supports arbitrary-shaped masks, facilitating more creative generation. (b) Additionally, it performs image inpainting with input images. Notably, our model demonstrates robust generalization, consistently producing ideal layouts across different seeds, with detailed experiments in the Appendix.
@@ 1 Introduction
Recently, diffusion models Rombach et al. (2022); Esser et al. (2024); Labs (2024) have significantly advanced text-to-image generation, enabling the creation of high-quality images through global textual descriptions. However, global text prompts alone are insufficient for precise image design control, as natural language lacks pixel-level precision, and image models often struggle to accurately model spatial relationships and object quantities Podell et al. (2024), significantly limiting their broader applicability. Consequently, integrating diverse conditions into text-to-image models has become a central research focus. For example, works like ControlNet Zhang et al. (2023) and IP-Adapter Ye et al. (2023) have introduced conditions such as depth maps, canny edges, and normal maps to influence the layout and style of generated images. Nevertheless, these conditions operate at a global image level, rather than targeting specific entities within the image. As a result, they lack the capability to provide fine-grained, entity-level control, which is essential for independently manipulating individual entities’ positions, shapes, or semantic attributes. This limitation impedes the generation of complex, highly customized images that require precise control over each entity.
In this paper, we focus on adding entity-level control for text-to-image models, a condition also known as instance-level or open-set layout control. This task allows for the input of multiple entity conditions, each including corresponding spatial and semantic controls. Previous researches Li et al. (2023); Zhou et al. (2024); Wang et al. (2024b); Wu et al. (2024) in this area have primarily relied on UNet-based Stable Diffusion models Podell et al. (2024). These methods typically encode entity spatial conditions in two ways: (1) representing positions via bounding boxes with Fourier encoding and auxiliary modules, or (2) mapping spatial information into attention maps to guide UNet’s cross-attention layers. However, these approaches have notable limitations. Firstly, bounding boxes restrict the model’s creativity and flexibility, as users often prefer to input arbitrary-shaped instance masks. Training with precise bounding box coordinates makes the network overly sensitive to numerical inputs, compromising the model’s robustness and leading to generated images with subpar aesthetic quality. Secondly, existing methods are UNet-based, while recent advancements in text-to-image generation have shifted to Diffusion Transformers (DiT) Esser et al. (2024); Labs (2024). This architectural gap makes previous designs obsolete and leaves entity-level control in DiT unexplored.
@@
Figure 2: Entity control ability of regional attention. The global prompt is “top-down view of a desk, laptop, a pot of rose, and a book.” (a) No entity control. (b) Untrained regional attention modifies regional details but lacks layout control. (c) After training, EliGen successfully achieves control over all entities.
@@
Unlike previous approaches that rely on bounding boxes, we aim to extend attention in DiT’s to achieve entity control with arbitrary-shaped spatial masks. To integrate entity controls into DiT, which accepts global prompt and latent image embeddings as inputs, we form the extended prompt sequence by combining local and global prompt embeddings. Furthermore, based on spatial conditions, we construct joint attention masks (entity-latent, inter-entity, and intra-entity). This enables regional attention for each entity through extended attention sequences and composed masks. Since local and global prompts share the same embedding space and no additional parameters are introduced, regional attention is effective even without training, as shown in Figure 2. In Figure 2b, the entity “yellow rose” successfully modifies regional details, while “laptop” and “book” show no noticeable changes, indicating that local prompts can influence generation but the train-free model lacks effective layout control abilities.
@@
Based on the above findings, we further consider training to enhance the model’s positional control capabilities. For the entity control task, each training sample requires an image, a global prompt, local prompts, and masks. We then construct a dataset of 500k high-quality annotated training samples using Flux to generate images and Qwen2-VL Wang et al. (2024a) vision-language model to annotate global prompts and entity information. Since no additional redundant structures are introduced, we adopt LoRA for efficient fine-tuning to ensure faster convergence.
@@
After training, our model, EliGen, acquires precise and robust entity-level control capabilities, seamlessly accepting arbitrary-shaped positional inputs, as shown in Figure 1a and 2c. Furthermore, we introduce a region-based noise fusion operation termed inpainting fusion, enabling EliGen to achieve impressive inpainting effects, as illustrated in Figure 1b. Leveraging the scalability of regional attention, EliGen can perform inpainting for multiple entities in a single forward pass. Additionally, EliGen can seamlessly integrate with other community-developed open-source models, such as IP-Adapter Team (2024) and MLLM Huang et al. (2024), to achieve higher levels of creativity.
We make the following contributions:
@@
  • We propose and finetune a regional attention mechanism to focus attention on specific image regions, enabling robust and precise entity-level control in EliGen.
  • Extensive experiments show that EliGen excels in entity-controlled generation, while performing effectively in image inpainting and seamless integration with community models, highlighting its creative potential.
@@ 2 Related Works
Text-to-Image Diffusion Models Ho et al. (2020); Song et al. (2021); Li et al. (2024) have catalyzed the evolution of generative AI by enabling the synthesis of high-quality images from textual descriptions. Latent Diffusion Models (LDMs) Rombach et al. (2022); Podell et al. (2024) have emerged as a pivotal innovation, operating in a compressed latent space rather than directly in pixel space, which significantly reduces computational costs while preserving synthesis quality. These models primarily incorporate text conditions through the cross attention module within the UNet architecture. The Diffusion Transformer (DiT) Peebles and Xie (2022) represents the first diffusion model to utilize a Transformer, integrating conditions into image latents via adaptive LayerNorm. SD3 Esser et al. (2024) and Flux Li et al. (2024) are the state-of-the-art models that adopt the DiT, enhancing image generation by concatenating text and latent image modalities and performing joint attention, thereby fully leveraging the semantic information of text to produce images of higher quality and finer detail. Despite the impressive capabilities of these models, generating images solely from text remains constrained. The community has been actively exploring methods to introduce additional conditions Zhang et al. (2023); Mou et al. (2023); Ye et al. (2023), with entity-level control being one such sophisticated condition.
@@
To equip foundational models with fine-grained control capabilities, researchers initially explored train-free methods such as diffusion fusion Bar-Tal et al. (2023) and cross attention guidance Ma et al. (2024). However, these approaches often produced unstable results and faced significant issues with entity omission. To tackle these challenges, researchers introduced additional modules to encode spatial information for each entity using bounding boxes and conducted task-specific supplementary training. GLIGEN Li et al. (2023) was the first entity control model to undergo training, integrating an additional gated attention module and encoding bounding box values through Fourier embedding. Subsequently, IFAdapter Wu et al. (2024) and MIGC Zhou et al. (2024) introduced new components: the instance feature adapter and the shading aggregation controller, respectively. They also mapped spatial information onto attention maps to guide cross attention processes. InstanceDiffusion Wang et al. (2024b) further refined the representation of spatial information for entities by categorizing it from coarse to fine, designing unique tokenizers for each category. While these training-based methods primarily focus on designing complex structures to encode spatial information, this paper shifts the focus to designing and training regional attention mechanisms without extra parameters. These mechanisms directly incorporate spatial information by controlling the regions where entity semantics exert influence.
@@ 3.1 Preliminaries
Here, we adopt the entity mask $m_{i}$ as a general representation for spatial localization, which is a more refined and flexible representation compared to the bounding box $b_{i}=\{x_{1},y_{1},x_{2},y_{2}\}$.
Figure 3: The regional attention mechanism within the double-stream transformer block of DiT. The global prompt is “A kitten sitting on a cloud”, and local entities are “kitten” and “cloud”. (a) The global and local prompts are encoded and then concatenated with the latent embeddings $z$ to form the sequence that requires attention. (b) The mask for regional attention is constructed from multiple components, each defining the specific region for which each sequence token should perform attention. In the composed mask $\mathrm{M}$, all colored regions indicate 1, and gray areas indicate 0.
@@
The core challenge lies in accurately encoding the local prompt $p_{i}$ and spatial localization $l_{i}$ of each entity and integrating them into the diffusion model. For $p_{i}$, this can be addressed similarly to the global prompt by extracting text embedding with a T5 text encoder Raffel et al. (2020). However, for $l_{i}$, prior methods often relied on hard-coding approaches to acquire numerical embedding, such as encoding bounding box coordinates via Fourier embedding or directly using mask images. Such hard-coded methods are suboptimal for the following reasons: (1) Rectangular Control Limitation: Encoding with bounding boxes restricts the model to rectangular control conditions, thereby losing the ability to handle arbitrary shapes, as illustrated in Figure 1a. (2) Sensitivity to Spatial Shapes: Hard-coded numerical values of bounding boxes or masks makes the model sensitive to the coordinates and shapes of entities, which reduces the model’s generalizability. (3) Additional Structural Requirements: Hard-coded approaches necessitate the design of extra structures to input localization embeddings.
We utilize a soft-coding scheme to encode $l_{i}$. Instead of using numerical spatial inputs, we interpret it as the region where local prompts $p_{i}$ exert their influence. Initially, we design a regional attention mechanism for DiT to handle each entity condition $e_{i}$. While untrained local prompts can semantically modify regions through regional attention, the model lacks the ability to precisely control layouts. To address this, we construct a dataset with entity annotations and finetune the model, significantly activating its ability to control layouts effectively.
@@ 3.2 Regional Attention
This work seeks to extend the attention mechanism to achieve refined control over multiple entities. We adopt the state-of-the-art DiT-based Flux Labs (2024) model as our baseline, while ensuring that the proposed regional attention mechanism can be easily applied to other text-to-image models. The core operation of DiT involves performing joint attention on the text and latent embedding. Given a text embedding $p$ with $n_{p}$ tokens encoded by the T5 text encoder and a latent embedding $z$ with $n_{z}$ tokens obtained from a noisy latent image through the patchify operation, DiT concatenates them as a single sequence and do Self-Attention jointly:
@@
The core mechanism of regional attention is illustrated in Figure 3. The illustration is based on the double-stream Transformer block. For the single-stream variant, the overall structure remains the same as the double-stream, except that the Text MLP and Latent MLP share weights. We explain this using an example involving $k=2$ entities. Firstly, as both the global prompt $p$ and the local prompt $p_{i}$ of entity $e_{i}$ are text, we use the T5 text encoder to extract their embeddings, with each prompt consisting of $n_{p}$ tokens. Subsequently, we concatenate them and use a Text MLP to obtain the attention input corresponding to the prompts, which is:
@@
Here, we construct the binary mask $\mathrm{M}$ to specify the regions that each token should attend to during attention, as shown in Figure 3b, which primarily includes four categories of regions. For consistency, we treat the global prompt as a global “entity”, with its entity mask being 1 for the entire image.
Entity-Latent: The latent embedding $z$ is derived by patchifying the latent image, where each token corresponds to a specific spatial patch within the image. Leveraging this spatial correspondence, we apply the same patchify operation to transform the entity mask $m_{i}$ into a vector that matches the length and positional alignment of $z$ in Eq. 9. By applying this vector to each token of the local prompt $p_{i}$ (comprising $n_{p}$ tokens in total), we can constrain $p_{i}$ to attend exclusively to the designated regions of $z$.
@@
After constructing the above regions, we need to concatenate them together to form a $N_{r}\times N_{r}$ composed attention map $\mathrm{M}$ as Figure 3b, where $N_{r}=(k+1)*n_{p}+n_{z}$. With $\mathrm{M}$, the regional attention mechanism defined in Eq. 10 can be expressed in detail as:
@@ 3.3 Dataset with Entity Annotation
Figure 4: The dataset generation and annotation pipeline.
For the entity control task, each training sample should consist of an image, a global prompt, and multiple entity annotations. Regarding the image source, prior works typically use open-source datasets like Laion LAION (2022) for images, but we observe that Laion’s image distribution differs significantly images generated by base model, leading to stylistic inconsistencies. Using Laion for training would distort the model distribution and degrade image quality (see Appendix), as the model would need to fit this mismatch. Therefore, we choose to directly use images generated by our baseline model, Flux, to avoid such distribution discrepancies.
The dataset generation pipeline is illustrated in Figure 4. Starting with DiffusionDB, we randomly select a caption and use the Flux model to generate the source image. Then, we employ Multimodal Large Language Mode (MLLM) Wang et al. (2024a) to re-caption the image, forming the global prompt. Unlike previous methods relying on Grounding DINO Liu et al. (2023), we directly use Qwen2-VL 72B, which possesses the most powerful grounding capabilities across all MLLMs, to randomly select entities in the image and annotate their local prompts and localization (via bounding boxes). This ensures consistency between local and global prompts. Since our regional attention employs soft-coded spatial localization, training with rectangular masks is sufficient to enable the model to generalize to arbitrary-shaped masks.

### v2

@@ Abstract
Recent advancements in diffusion models have significantly advanced text-to-image generation, yet global text prompts alone remain insufficient for achieving fine-grained control over individual entities within an image. To address this limitation, we present EliGen, a novel framework for Entity-Level controlled Image Generation. We introduce regional attention, a mechanism for diffusion transformers that requires no additional parameters, seamlessly integrating entity prompts and arbitrary-shaped spatial masks. By contributing a high-quality dataset with fine-grained spatial and semantic entity-level annotations, we train EliGen to achieve robust and accurate entity-level manipulation, surpassing existing methods in both spatial precision and image quality. Additionally, we propose an inpainting fusion pipeline, extending EliGen’s capabilities to multi-entity image inpainting tasks. We further demonstrate EliGen’s flexibility by integrating it with other open-source models such as IP-Adapter, In-Context LoRA and MLLM, unlocking new creative possibilities. The source code, model, and dataset are published at the Project Page.
@@ 1 Introduction
Recently, diffusion models Rombach et al. (2022a); Esser et al. (2024); Labs (2024) have significantly advanced text-to-image generation, enabling the creation of high-quality images through textual descriptions. However, a single prompt is insufficient for precise image design control, as natural language cannot describe image content with pixel-level precision, and image models often struggle to accurately model spatial relationships and object quantities Podell et al. (2024). This pitfall significantly limits their broader applicability. Consequently, integrating diverse conditions into text-to-image models has become a central research focus. For example, recent studies like ControlNet Zhang et al. (2023) and IP-Adapter Ye et al. (2023) have introduced conditions such as depth maps, normal maps and reference images to control the layout and style of generated images. Nevertheless, these conditions operate at a global image level, rather than targeting specific entities within the image. As a result, they lack the capability to provide fine-grained and entity-level control. This limitation impedes the generation of complex, highly customized images, motivating us to investigate precise controlling approaches for entity-level control.
@@
Figure 1: Entity control ability of EliGen. The global prompt is “top-down view of a desk, laptop, a pot of rose, and a book.” (a) No entity control. (b) Untrained regional attention modifies regional details (entity “rose”) but lacks position control ability (entities “laptop” and “book”). (c) After training, EliGen successfully achieves control over all entities.
Figure 2: EliGen enables spatial and semantic control of each entity. (a) By incorporating local prompts and masks for each entity, it generates images with specific layouts and details. Unlike previous models restricted to rectangular controls, EliGen supports arbitrary-shaped masks, facilitating more creative generation. (b) Additionally, it performs image inpainting with input images. Notably, our model demonstrates robust generalization, consistently producing ideal layouts across different seeds, with detailed experiments in the Supplementary Material.
In this paper, we focus on adding entity-level control for text-to-image models. Entity-level control is also known as instance-level or open-set layout control. This task allows for the input of multiple entity conditions, each including corresponding spatial and semantic information. Previous studies Li et al. (2023); Zhou et al. (2024); Wang et al. (2024b); Wu et al. (2024) have preliminarily demonstrated the feasibility of entity-level control based on the Stable Diffusion series models Rombach et al. (2022b); Podell et al. (2024). These methods typically encode spatial conditions in two ways: (1) representing positions via bounding boxes with Fourier encoding Mildenhall et al. (2021) and auxiliary modules, or (2) mapping spatial information into attention maps of cross-attention layers. However, these approaches have notable limitations. Firstly, bounding boxes restrict the model’s flexibility, as users prefer to input arbitrary-shaped entity masks. Training with precise bounding box coordinates makes the network overly sensitive to numerical inputs, compromising the model’s robustness and leading to generated images with subpar aesthetic quality. Secondly, existing methods are closely coupled with the model structure of UNet Ronneberger et al. (2015), while recent advancements in text-to-image generation have proved that Diffusion Transformers (DiT) Esser et al. (2024); Labs (2024) are capable of achieving superior performance. Previous methods lack generality and are unable to adapt to the evolving text-to-image models.
@@
Unlike previous approaches that rely on bounding boxes, we aim to generally extend the attention modules of diffusion models to achieve entity-level control with arbitrary-shaped spatial information. To integrate entity conditions into the attention modules, we employ structured text embeddings to represent semantic information, which includes a global prompt describing the overall content of the image and several local prompts describing specific details. Each local prompt is accompanied by a mask representing the corresponding region. Furthermore, based on spatial conditions, we construct joint attention masks (entity-latent, inter-entity, and intra-entity). This enables regional attention for each entity through the extended attention sequences and composed masks. Since local and global prompts share the same embedding space and no additional parameters are introduced, the train-free regional attention mechanism is able to modify entity details. However, it still lacks precise layout control capabilities, as shown in Figure 1.
@@
Based on the above findings, we further consider training to enhance the model’s entity-level control capabilities. For this task, each training sample requires an image accompanied by a global prompt, several local prompts, and corresponding masks. We construct a dataset consisting of 500k high-quality annotated training samples using FLUX.1-dev for image generation and Qwen2-VL Wang et al. (2024a) for image captioning and entity annotation. Since no additional structures are introduced, we adopt LoRA Hu et al. (2021) for efficient fine-tuning to ensure faster convergence.
@@
After training, our model, EliGen, acquires precise and robust entity-level control capabilities, seamlessly accepting arbitrary-shaped positional inputs, as shown in Figure 2a and 1c. Qualitative and quantitative experiments show that EliGen outperforms previous methods in both spatial accuracy and image quality. Furthermore, we introduce a region-based noise fusion operation termed inpainting fusion, enabling EliGen to achieve impressive inpainting effects, as illustrated in Figure 2b. Additionally, EliGen can seamlessly integrate with other open-source models, such as IP-Adapter Team (2024), In-Context LoRA Huang et al. (2024a), and MLLM Huang et al. (2024b), to achieve higher levels of creativity, such as styled entity control, entity transfer, and dialogue-based image design and editing.
The contributions of this paper are summarized as follows:
@@
  • We present a regional attention mechanism and fine-tune it to focus attention of entites on specific regions, enabling precise entity-level control in EliGen.
  • Extensive experiments show that EliGen excels in entity-controlled generation. It also performs effectively in image inpainting and seamlessly integrates with community models, proving its strong creative potential.
@@ 2 Related Work
Text-to-Image Diffusion Models Ho et al. (2020); Song et al. (2021); Li et al. (2024) have catalyzed the evolution of generative AI by generating high-quality images from textual descriptions. Latent Diffusion Models (LDMs) Rombach et al. (2022a); Podell et al. (2024) have emerged as a pivotal innovation, operating in a compressed latent space rather than directly in pixel space, which significantly reduces computational costs. These models primarily incorporate text conditions through the cross attention module within the UNet Ronneberger et al. (2015) architecture. The Diffusion Transformer (DiT) Peebles and Xie (2022) represents the first diffusion model to utilize a Transformer, integrating conditions into image latents via adaptive LayerNorm. SD3 Esser et al. (2024) and FLUX Labs (2024) are the state-of-the-art models that adopt the DiT, enhancing image generation by concatenating text and latent image modalities, thereby fully leveraging the semantic information of text to produce images of higher quality and finer detail. Despite the impressive capabilities of these models, generating images solely from text remains constrained. Recent studies have been exploring methods to introduce additional conditions Zhang et al. (2023); Mou et al. (2023); Ye et al. (2023), with entity-level control being one such sophisticated condition.
@@
To equip foundational models with fine-grained control capabilities, researchers originally explored train-free methods such as diffusion fusion Bar-Tal et al. (2023) and cross attention guidance Ma et al. (2024). However, these approaches often produce unstable results and encounter entity omission issues. To tackle these challenges, researchers introduced additional modules to encode spatial information for each entity using bounding boxes and conducted task-specific training. GLIGEN Li et al. (2023) is the first entity control model to undergo training, integrating an additional gated attention module and encoding bounding boxes through Fourier embedding. Subsequently, MIGC Zhou et al. (2024) introduces the shading aggregation controller and maps spatial information onto attention maps to guide cross attention processes. InstanceDiffusion Wang et al. (2024b) further categorizes spatial information from coarse to fine and designs unique tokenizers for each category. While these studies focus on designing complex structures to encode spatial information, this paper shifts the focus to designing and training regional attention mechanism without extra parameters and incorporating spatial information by controlling the regions where entity semantics exert influence.
@@ 3.1 Preliminaries
Here, we adopt the entity mask $m_{i}\in\{0,1\}^{h\times w}$ as a general representation for spatial localization, which is a more refined and flexible representation compared to the bounding box $b_{i}=\{x_{1},y_{1},x_{2},y_{2}\}$.
Figure 3: The regional attention mechanism within the double-stream transformer block of DiT. (a) The diffusion model. (b) The global and local prompts are encoded and concatenated with the latent embeddings $z$ to form the attention sequence. (c) The attention mask $\mathrm{M}$ is constructed from multiple components, each defining the specific region for which each sequence token should perform attention. In the composed mask $\mathrm{M}$, all colored regions indicate 1, and gray areas indicate 0.
@@
The core challenge lies in accurately encoding the local prompt $p_{i}$ and spatial localization $l_{i}$ of each entity and integrating them into the diffusion model. For $p_{i}$, integration can be addressed similarly to the global prompt by extracting text embedding with a T5 text encoder Raffel et al. (2020). However, for $l_{i}$, prior methods often relied on hard-coding approaches to acquire numerical embedding, such as encoding bounding box coordinates via Fourier embedding or directly using mask images. Such hard-coded methods are suboptimal for the following reasons: (1) Rectangular Control Limitation: Encoding with bounding boxes restricts the model to rectangular control conditions, thereby losing the ability to handle arbitrary shapes, as illustrated in Figure 2a. (2) Sensitivity to Spatial Shapes: Hard-coded numerical values of bounding boxes or masks makes the model sensitive to the coordinates and shapes of entities, which reduces the model’s generalizability. (3) Additional Structural Requirements: Hard-coded approaches necessitate the design of extra structures to input localization embeddings.
We utilize a soft-coding scheme to encode $l_{i}$. Instead of using numerical spatial inputs, we interpret it as the region where local prompts $p_{i}$ exert their influence. Initially, we design a Regional Attention (RA) mechanism to handle each entity condition $e_{i}$. While untrained RA can semantically modify regional details, it lacks the ability to precisely control layouts. To address this, we construct a dataset with entity annotations and finetune the model, significantly activating its ability to control layouts effectively.
@@ 3.2 Regional Attention
This work seeks to extend the attention mechanism to achieve refined control over multiple entities. We adopt the state-of-the-art DiT-based FLUX.1-dev Labs (2024) model as our baseline, while ensuring that the proposed regional attention mechanism can be easily applied to other text-to-image models. The core operation of DiT involves performing joint attention on the text and latent embedding. Given a text embedding $p$ with $n_{p}$ tokens encoded by the T5 text encoder and a latent embedding $z$ with $n_{z}$ tokens obtained from a noisy latent image through the patchify operation, DiT concatenates them as a single sequence and do Self-Attention jointly:
@@
The core mechanism of regional attention is illustrated in Figure 3. The illustration is based on the double-stream transformer block. For the single-stream variant, the overall structure remains the same, except that the Text MLP and Latent MLP share weights. We explain this using an example involving $k=2$ entities. Firstly, as both the global prompt $p$ and the local prompt $p_{i}$ of entity $e_{i}$ are text, we use the T5 text encoder to extract their embeddings, with each prompt consisting of $n_{p}$ tokens. Subsequently, we concatenate them and use a Text MLP to obtain the attention input corresponding to the prompts, which is:
@@
Here, we construct the binary mask $\mathrm{M}$ to specify the regions that each token should attend to during attention, as shown in Figure 3c, which primarily includes four categories of regions. For consistency, we treat the global prompt as a global “entity”, with its entity mask being 1 for the entire image.
Entity-Latent: The latent embedding $z$ is derived by patchifying the latent image, where each token corresponds to a specific spatial patch within the image. Leveraging this spatial correspondence, we apply the same patchify operation to transform the entity mask $m_{i}$ into a vector that matches the length and positional alignment of $z$ in Eq. 9. By applying this vector to each token of the local prompt $p_{i}$, we can constrain $p_{i}$ to attend exclusively to the designated regions of $z$.
@@
After constructing the above regions, we need to concatenate them together to form a $N_{r}\times N_{r}$ attention map $\mathrm{M}$ as Figure 3c, where $N_{r}=(k+1)*n_{p}+n_{z}$. With $\mathrm{M}$, the regional attention mechanism defined in Eq. 10 can be expressed in detail as:
@@ 3.3 Dataset with Entity Annotation
For the entity control task, each training sample should consist of an image, a global prompt, and multiple entity annotations. Regarding the image source, prior works typically use open-source datasets like Laion LAION (2022) for images, but we observe that Laion’s image distribution differs significantly images generated by base model, leading to stylistic inconsistencies. Using Laion for training would distort the model distribution and degrade image quality (see the Supplementary Material for ablations), as the model would need to fit this mismatch. Therefore, we choose to directly use images generated by our baseline model, FLUX.1-dev, to avoid such distribution discrepancies.
The dataset generation pipeline and statistics are illustrated in the Supplementary Material. Starting with DiffusionDB, we randomly select a caption and use FLUX.1-dev to generate the source image. Then, we employ Multimodal Large Language Mode (MLLM) Wang et al. (2024a) to re-caption the image, forming the global prompt. Unlike previous methods relying on Grounding DINO Liu et al. (2023), we directly use Qwen2-VL 72B, which possesses the most powerful grounding capabilities across all MLLMs, to randomly select entities in the image and annotate their local prompts and localization (via rectangular masks). This ensures consistency between local and global prompts. Since our regional attention employs soft-coded spatial localization, training with rectangular masks is sufficient to enable the model to generalize to arbitrary-shaped masks.